Please help me solve this. I am so lost it's pathetic.
x²+9=34

Answer by checkley77(12844)

x^2+9=34
x^2+9-34=0
x^2-25=0
(x+5)(x-5)=0
x+5=0
x=-5 ans.
x-5=0
x=5 ans.
Proof:
5^2+9=34
25+9=34
34=34